40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 20 20:20:36 2020, gautam, Update, PSL, PMC servo checkout PMCsweep.pdfOLTFmeas.pdf
    Reply  Wed Jan 22 11:22:39 2020, gautam, Update, PSL, PMC modulation depth measurement modDepth.pdf
       Reply  Wed Jan 22 20:12:36 2020, gautam, Update, PSL, PMC demodulator electrical characterization demodChar.pdfmixerChar.pdf
          Reply  Thu Jan 23 14:37:05 2020, gautam, Update, PSL, PMC VGA chip damaged? VGAchar.pdf
             Reply  Thu Jan 23 16:37:14 2020, rana, Update, PSL, PMC VGA chip damaged? 
             Reply  Thu Jan 23 22:10:01 2020, gautam, Update, PSL, PMC servo pulled out 
                Reply  Fri Jan 24 15:42:08 2020, gautam, Update, PSL, PMC servo restored 
                   Reply  Sun Jan 26 13:47:00 2020, gautam, Update, PSL, PMC servo characterization elecTFs.pdfVGAchar_postFix.pdfVGAlinearity_postFix.pdfnewOLTFs.pdf
                      Reply  Tue Jan 28 14:33:24 2020, gautam, Update, PSL, Inferred free-running frequency noise inLoopNoise_IMClocked.pdffreqNoiseComparison.pdf
          Reply  Thu Jan 23 18:52:31 2020, gautam, Update, PSL, PMC RFPD characterization PDresp.pdf
Message ID: 15135     Entry time: Mon Jan 20 20:20:36 2020     Reply to this: 15139
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC servo checkout 

Summary:

The PDH discriminant of the PMC servo was measured to be ~0.064 GV/m. This is ~50 times lower than what is reported here. Perhaps this is a signature of the infamous ERA decay, needs more investigation.

Details:

  • Calibration of the error and control points were done using 1 Hz triangle wave injection to the "EXT DC" input of the PMC servo. Two such sweeps are shown in Attachment #1 (measured data as points, fits as solid lines). For the control signal monitor, I've multiplied the signal obtained on the scope by 49.6, which is the voltage divider implemented for this monitor point.
  • The PDH discrimiannt was calibrated into physical units knowing the modulation frequency of the PMC, which is 35.5 MHz. The error in this technique due to the free-running NPRO frequency noise is expected to be small since the entire fringe is crossed in <30 ms, in which time the laser frequency is expected to change by < 5 kHz.
  • The drive to the PZT was calibrated into physical units using the same technique. This number is within a factor of 2 of the number reported here
  • Attachment #2 shows the loop OLTF measured using the usual IN1/IN2 prescription (with an SR560). In fact, the 8kHz feature makes the loop unstable. For convenience, I've overlaid the OLTF from March 2017, when things were running smoothly. It is not clear to me why even though the optical gain is now lower, a smaller servo gain results in a larger UGF.

The light level hasn't changed by a factor of 50, leading me to suspect the modulation depth. Recall that the demodulation of the PMC is now done off the servo board using a minicircuits mixer (hence, the "C1:PSL-PMC_LODET" channel isn't a reliable readback of the LO signal strength over time). Although there is a C1:PSL-PMC_MODET channel which looks like it comes from the crystal reference card, and so should still work - this, however, shows no degradation over 1 year.

Somebody had removed the BLP-1.9 that I installed at the I/F output of the mixer to remove the sum frequency component in the demodulated signal, I reinstalled this. I find that there are oscillations in the error signal if the PMC servo gain is increased above 14.5 on the MEDM slider.

Attachment 1: PMCsweep.pdf  90 kB  Uploaded Tue Jan 21 10:24:01 2020  | Hide | Hide all
PMCsweep.pdf
Attachment 2: OLTFmeas.pdf  155 kB  Uploaded Tue Jan 21 11:54:09 2020  | Hide | Hide all
OLTFmeas.pdf
ELOG V3.1.3-